OpenStudy (samanthagreer):

Simplify: 6m + 7n + 5m – 3n A. 15mn B. 11m + 4n C. 11m + 10n D. m + 10n i think b

OpenStudy (zaynab123):

yes thats correct

OpenStudy (albert0898):

Combine like terms 11m + 4n

OpenStudy (zaynab123):

6m+4m=11m and 7n-3n=4n

OpenStudy (zaynab123):

i meant 6+5=11

OpenStudy (samanthagreer):

thanks guys

OpenStudy (alivejeremy):

Yw